OpenStudy (henrietepurina):

Algebra 1 question Fan + Medal

OpenStudy (henrietepurina):

Based on the graph below, what are the solutions to the equation f(x) = g(x)? graph of function f of x equals 2 times x plus 5. Graph of function g of x equals x cubed minus 3 times x squared minus 4 times x plus 12

OpenStudy (anonymous):

f(x)=g(x) 2x+5=x^3-3x^2-4x+12 f(x)=g(x) is at the points where the two curves intersect.

OpenStudy (henrietepurina):

could it be x=8 @surjithayer

OpenStudy (henrietepurina):

@nikato

OpenStudy (bohotness):

i can help

OpenStudy (henrietepurina):

thanks :)

OpenStudy (henrietepurina):

could the answer be x=8 @bohotness ?

OpenStudy (bohotness):

(0,2)

OpenStudy (bohotness):

and (-2, -12)

OpenStudy (bohotness):

it's D

OpenStudy (bohotness):

because those are all the x-axis values

OpenStudy (henrietepurina):

that is not one of the answer choices :O

OpenStudy (henrietepurina):

x = -3, 0.8, 4 x = -1.9, 0.8, 4.1 x = -2, 2, 3 x = -2.5, 0.8, 3

OpenStudy (bohotness):

o./o

OpenStudy (henrietepurina):

ikr

OpenStudy (henrietepurina):

hmm... let me check for any writing mistakes

OpenStudy (henrietepurina):

nope.. nothing :O

OpenStudy (bohotness):

o.o

OpenStudy (henrietepurina):

mind trying again.... just incase?

OpenStudy (henrietepurina):

I am lenient towards D... you?

OpenStudy (anonymous):

from the graph it appears x=-1.9,0.8,4.1